OpenStudy (beccab003):

Can someone check my answer? Thanks! Find f(-2): F(x) = 2x^3 – 3x/x^2 + 1 4.4 2 -2 -4.4 It is -2. Medal to best answer. Thanks!

jhonyy9 (jhonyy9):

f(-2) mean thet just substitute the -2 in place of x right ?

OpenStudy (beccab003):

Yep

OpenStudy (beccab003):

That is what I did.

jhonyy9 (jhonyy9):

so f(-2)= (2*(-2)^3 -3*(-2))/((-2)^2 +1) = do you can calculi it ?

jhonyy9 (jhonyy9):

f(-2)=( 2*(-8) +6)/(4+1) = -10/5 = - 2

OpenStudy (beccab003):

THANKS